Rewrite as proper complex number (a+bi)

(5/8+2i)/(1/2-1/4i)
1
Expert's answer
2010-11-22T14:28:07-0500
(5/8+2i)/(1/2-1/4i) = 5/16 + i & - 5/32 i & +1/2 & = 13/16 + 27/32 i